there are two charges, -3.55 microC at the origin, and +4.40 microC at the point (x,y)= (2.63m,3.60m). What is the magnitude and direction of the electrostatic force acting on the +4.38microC charge?
Answer #1

Q1 = - 3.55 x 10-6 C; Q2 = 4.40 x 10-6 C

Distance between the charges, r = (2.632 + 3.602 )1/2 = 4.46 m

Electric force acting on the charge Q2 is F = 1/4πε0 Q1Q2/r2

F = 9 x 109 x 3.55 x 10-6 x 4.40 x 10-6 /4.462 =7.07 x 10-3 N

The magnitude of the electrostatic force on the +4.38microC charge is F = 7.07 x 10-3 N and this force is directed towards -3.55 micro C charge.